OpenStudy (anonymous):

Solve for x in the proportion 3x over 3x - 36 = -1 over x

OpenStudy (anonymous):

3x / ( 3x -36) = -1/x ?

OpenStudy (anonymous):

yea

jhonyy9 (jhonyy9):

3x/3(x-12) = -1/x x/(x-12) = -1/x cross multiplie x^2 = - (x-12) x^2 = -x+12 x^2 +x -12 =0 (x+4)(x-3)=0 x+4=0 x_1=-4 and x-3=0 x_2=3 x_1 and x_2 are solutions

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

k

jhonyy9 (jhonyy9):

good luck bye